evaluate (1+i)^12 by using De Moivre's Theorem.express the result in rectangular form
Answer by Alan3354(30993)

z = 1+i
----------
(1+i)^12
r = |i+i| = sqrt(2)
theta = 45º
z^12 = r^12cis(theta*12)
= 64cis540
= 64*(cos(540 + isin(540))
= 64*(-1 + i*0)
= -64